OpenStudy (anonymous):

hose A can fill a certain vat in 3 hours. after 2 hours of pumping, hose A it tuned off. hose B is then the tuned on and completes filling the vat in 3 more hours. how long would it take hose B to fill the vat alone?.

OpenStudy (anonymous):

hose A can fill a certain vat in 3 hours. after 2 hours of pumping, hose A it tuned off. hose B is then the tuned on and completes filling the vat in 3 more hours. how long would it take hose B to fill the vat alone?.

OpenStudy (barrycarter):

Hint: what portion of the vat can Hose A fill per hour?

OpenStudy (cruffo):

If A ran for 2 hours, then the vat is 2/3 full. B only need to fill the remaining 1/3, which it does in 3 hours. So the equation to solve is \[\Large \frac{2}{3} + \frac{3}{x} = 1\]

OpenStudy (cruffo):

(assuming a constant rate from both hoses :)

OpenStudy (anonymous):

If B can fill 1/3 tank in 3 hours, then B can fill 1 tank in x hours. Solve the following for x :\[\frac{1/3}{3}=\frac{1}{x} \]x=9 hours